A car loan requires 60 monthly payments of $250 and has a 6% APR.1 (a) What is the present value of the loan one month before the first payment is due? 


1
Expert's answer
2020-09-16T20:01:06-0400

Interest rate per annum = 6% (6%/12 = 0.005 per month)

Total payments: $250.25

Payment (PMT) = $250 per month

n = 60 monthly payments

Future value (FV)

"FV = PMT * \frac{(1-(1+r)^{-n}}{r}"

"FV = 250 * \frac{1-(1+0.06)^{-60}}{6\%/12}"

"FV = 250 * \frac{1-(1.06)^{-60}}{0.005}"

"FV = 250 * \frac{(1-0.030314)}{0.005}"

"FV=250 * \frac{0.969686}{0.005}"

FV = 48,484.30

Present value of the loan one month before the first payment is due

PV = FV / (1+r) n

PV = "48,484.30 \/ (1+0.005)^{60}"

PV = "48,484.30 \/ (1.005)^{60}"

PV = 48,484.30 / 1.34885

PV = $ 35,944.92